OpenStudy (goformit100):

Use the properties of sets to prove that for all the sets A and B A – (A∩B) = A – B

OpenStudy (goformit100):

@satellite73 Sir Help Me Please.

mathslover (mathslover):

I think, drawing venn diagram for it will work.

mathslover (mathslover):

Use De Morgan's law : \(A - (B \cap C ) = (A-B) \cup (A-C) \) We have : \(A - (A \cap B) = (A-A) \cup (A-B) \) \(0 \cup (A-B) = A-B \)

mathslover (mathslover):

Not zero, \(\phi \cup (A-B) = A-B\)

mathslover (mathslover):

Got it? @goformit100

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

Given 2 sets \(\bf A=\left\{ 1,2,3 \right\}\) and \(\bf B=\left\{ 2,3,4 \right\}\), we find that:\[\bf A ∩ B =\left\{ 1,2,3 \right\} ∩\left\{ 2,3,4 \right\}=\left\{ 2,3 \right\}\]\[\bf A-(A∩B)=\left\{ 1,2,3 \right\}-\left\{ 2,3 \right\}=\left\{ 1 \right\}\]\[\bf A-B=\left\{ 1,2,3 \right\}-\left\{ 2,3,4 \right\}=\left\{ 1 \right\}\]This concludes that:\[\bf A-(A∩B)=A-B\]But we must prove this equality for any general case for two given sets A and B for which you can do what mathslover stated.